(a) Find the inverse of 19 modulo 141, using the Extended Euclidean Algorithm.




Show your steps.

1
Expert's answer
2021-11-02T14:08:35-0400

Euclidean Algorithm:

"141=7\cdot 19+8"

"19=2\cdot 8+3"

"8=2\cdot 3+2"

"3=1\cdot 2+1"


We have:

"1= \mathbf{3}-\mathbf{2} \\ \ \ \ = \mathbf{3}-(\mathbf{8}-2\cdot \mathbf{3)}
\\ \ \ \ =3\cdot \mathbf{3}-\mathbf{8}
\\ \ \ \ =3\cdot (\mathbf{19}-2\cdot \mathbf{8})-\mathbf{8}
\\ \ \ \ =3\cdot \mathbf{19}-7\cdot \mathbf{8}
\\ \ \ \ =3\cdot \mathbf{19}-7\cdot (\mathbf{141}-7\cdot \mathbf{19})
\\ \ \ \ =52\cdot \mathbf{19}-7\cdot \mathbf{141}"


"1=52\cdot 19-7\cdot 141\equiv 52\cdot 19\mod 141"


Answer: "52" .